jhonyy9 (jhonyy9):

- can being this proof accepted - 1. substitution - let p and k , two prime numbers greater or equal 2,from the set of prime numbers, P, in the form : p=2a + 1 and k=2b + 1 , such that a and b are natural numbers,from the set of natural numbers N, - let m=2n ,m greater or equal 4,even number,from the set of natural numbers N and n grater or equal 2,natural number from set of natural numbers N, 2. conclusion - every even integer greater than 2 can be expressed as the sum of two primes 3. demonstration * - step 0. for n=2 --- m=4 --- 4=2+2 ** - step 1. for - if n is greater or equal

jhonyy9 (jhonyy9):

1. substitution - let p and k , two prime numbers greater or equal 2,from the set of prime numbers, P, in the form : p=2a + 1 and k=2b + 1 , such that a and b are natural numbers,from the set of natural numbers N, - let m=2n ,m greater or equal 4,even number,from the set of natural numbers N and n grater or equal 2,natural number from set of natural numbers N, 2. conclusion - every even integer greater than 2 can be expressed as the sum of two primes 3. demonstration * - step 0. for n=2 --- m=4 --- 4=2+2 ** - step 1. for - if n is greater or equal 3 so always will be a number a and b such that n=a + b + 1 - dem. 3=1+1+1 4=2+1+1 5=2 +2+1 ................ n=a+b+ 1 - so for n=k than k=a+b+1 - suppose that is true - for k+1=(a+b+1)+1=(k)+1=k+1 - so for k+1 is true - for n grater than 2 always will be a number a and b such that n =a+b+ 1 *** - step 2. - every even integer greater than 4 can be expressed as the sum of two primes m=p+k - demonstration by ,,reductio ad absurdum" - so than m is not equal p+k - so than 2n is not equal 2a+1+2b+1 - so than 2n is not equal 2a+2b+2 / divide both sides by 2 - so than n is not equal a+b+1 - so but n=a+b+1 was proved that is true - so than m=p+k is proved that is true

jhonyy9 (jhonyy9):

- here is the website code

jhonyy9 (jhonyy9):

where is publiched too

jhonyy9 (jhonyy9):

HA

jhonyy9 (jhonyy9):

HAPPY NEW YEAR !!!

jhonyy9 (jhonyy9):

<iframe src=" https://docs.google.com/present/embed?id=ddd7shbj_1hnwt4pcg&interval=5&autoStart=true&loop=true " frameborder="0" width="410" height="342"></iframe>

OpenStudy (anonymous):

good work guys